# Adaptive Layerwise Perturbation: Unifying Off-Policy Corrections for LLM RL

### 19 Mar 2026 v1

@@ 4 Experiments
We conduct experiments on single-turn reasoning and multi-turn agentic tasks respectively.
@@ 4.1.1 Experimental settings
In the single-turn reasoning, we study mathematical reasoning tasks without tool-calling.
Dataset.
We construct the training dataset by merging the math subset from Guru RL-92k [^3] https://huggingface.co/datasets/LLM360/guru-RL-92k(Cheng et al., 2025) and $75k$ OpenR1444https://huggingface.co/datasets/weqweasdas/from_default_filtered_openr1_with_scores(Hugging Face, 2025) subset. Guru contains $91.9K$ high-quality samples spanning six diverse reasoning-intensive domains, processed through a comprehensive five-stage curation pipeline to ensure both domain diversity and reward verifiability. OpenR1 consists of 220k math problems with two to four reasoning traces generated by DeepSeek R1 for problems from NuminaMath 1.5. We employ the Math-Verify tool (Kydlíček, ) for automatic solution correctness verification. The details are deferred to Appendix D.
Training.
The implementations are based on the verl framework (Sheng et al., 2024), and we follow most of the parameter settings in verl. Detailedly, we apply the AdamW optimizer with learning rate $1\times 10^{-6}$ and adopt the clip higher trick (Yu et al., 2025) to clip the sampling ratio to an asymmetric range $(1-\epsilon_{l},1+\epsilon_{h})$. We conduct Seq-ALP (sequence-level ALP Equation (3)) and Token-ALP (token-level ALP Equation (2)) under the same training pipeline. The training is initialized from Qwen2.5-Math-1.5B-base (Yang et al., 2024). In each iteration, we sample $512$ prompts and rollout $n=8$ samples with temperature $1.0$ for each prompt, and the policy update number is $16$ in each iteration.
Benchmarks.
We evaluate the models’ performance on $5$ benchmarks: Math500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), Olympiad Bench (He et al., 2024), AIME2024555https://huggingface.co/datasets/math-ai/aime24 and AIME2025666https://huggingface.co/datasets/math-ai/aime25. We use average@$32$ for evaluation, i.e., the accuracy is averaged over $32$ responses per prompt under temperature $1.0$. The models are allowed to generate $4096$ tokens. We compare with four baselines: Seq-Bypass, Seq-MIS, Token-MIS and vanilla GRPO that differ in how training-inference importance ratio $\pi_{\theta_{\mathrm{old}}}^{\mathrm{train}}/\pi_{\theta_{\mathrm{old}}}^{\mathrm{infer}}$ and policy update importance ratio $\pi_{\theta}^{\mathrm{train}}/\pi_{\theta_{\mathrm{old}}}^{\mathrm{train}}$ are formed and clipped. As provided in Table 1, in each row, token/seq means the aggregation way of the ratio is in token-level or sequence level. The definitions of those notations are provided in Sec. 3.1.
GRPO token None None
Token-MIS token seq None
Seq-MIS seq seq None
Seq-Bypass None None seq
Table 1: Settings of parameters of importance ratios
@@ 4.1.2 Main Results
Figure 3: Single-turn training dynamics under training-inference mismatch. We report step-wise (policy-update steps) optimization diagnostics (batch-averaged reward, gradient norm, policy entropy, and KL divergence between rollout and training policies) for GRPO, MIS, Bypass, and ALP. The reward mean is smoothed with a 10-step moving average.
ALP-type delivers the top performance.
Under single-turn evaluation in Table 2 across five math benchmarks, Token-ALP attains the highest overall average score ($37.87$), with Seq-ALP ranking second ($36.83$). Token-ALP leads on four out of five benchmarks (Math500, Minerva Math, Olympiad Bench, and AIME24), while Seq-ALP achieves the best result on AIME25 ($11.98$). In comparison, Token-MIS is the strongest baseline, while Seq-MIS and GRPO are slightly weaker overall ($35.54$ and $35.77$). Seq-Bypass performs the worst (Average $34.82$).
ALP consistently improves training stability.
The training dynamics in Figure 3 further explain these performance differences under training-inference mismatch. First, both token-level and seq-level ALP exhibit stable optimization throughout training: the gradient norm remains well-controlled (no sharp spikes), the policy entropy stays in a reasonable band (avoiding both collapse and runaway growth), and critically the training-inference KL remains near zero, indicating minimal drift between the rollout policy used for data collection and the policy effectively optimized during training. Second, methods without perturbation show clear instability signatures. In particular, GRPO and Seq-MIS undergo abrupt transitions late in training, characterized by large gradient-norm spikes, a pronounced training-inference KL surge, and a simultaneous entropy anomaly, which coincide with a sharp drop in reward mean – a hallmark of training collapse under mismatch. By contrast, Seq-Bypass and Token-MIS appear more stable than GRPO/Seq-MIS, but they still fall short of ALP in combining strong stability with the best overall accuracy: Token-MIS remains the strongest baseline, yet both Token-ALP and Seq-ALP provide consistently better aggregate performance while preserving stable training dynamics.
@@ 5.1 Experimental settings
We choose multi-turn Tool-Integrated Reasoning (TIR) task with a Python interpreter to perform mathematical reasoning, where the agent can code to validate results, solve equations and so on. We adapt the codebase from SimpleTIR (Xue et al., 2025) and mask void turns (turns without tool calls and without final answers) in the loss. For the dataset, we merge the dataset Math3-5 from SimpleRL (Zeng et al., 2025), Deepscaler (Luo et al., 2025), and rStar2-Agent (Shang et al., 2025). ALP follows zero-RL setting and starts from Qwen2.5-7B-base. During training, the rollout batch size is $512$, and the mini update size is $32$. The maximum response length is set as 1616K, with a maximum of $5$ turns of code execution. The baselines and test benchmarks are the same with the single turn setting.
@@ 5.2 Main Results
Figure 4: Training dynamics: The leftmost plot illustrates that Seq-ALP maintains higher entropy than MIS baselines, ensuring sufficient exploration, while avoiding the unstable entropy growth ("blow-up") observed in Seq-Bypass. The center and right plots show that Seq-ALP maintains consistent convergence in both KL metrics. In contrast, Token-MIS exhibits significant instability in Train-Inference KL, and Seq-MIS suffers from sharp spikes in Policy Update KL, highlighting the robustness of the ALP method.
Figure 5: Pass@k performance analysis on AIME 2024 and AIME 2025 datasets for TIR tasks. The results compare ALP against baseline methods across varying rollout numbers ($k$). ALP consistently achieves the highest accuracy in the range of $k=16\sim 256$, indicating that adaptive latent perturbation significantly enhances the model’s exploration efficiency and solution diversity.
ALP has the best performance.
On multi-turn TIR benchmarks in Table 3, Seq-ALP achieves the highest overall average accuracy ($50.53$), improving over Token-MIS ($48.74$), Seq-MIS ($46.94$), and Seq-Bypass ($46.66$). The gains are broad-based: Seq-ALP is best on Math500 ($84.29$), Minerva Math ($43.10$), Olympiad Bench ($52.75$), and AIME24 ($43.85$), and is competitive on AIME25 ($28.65$) (slightly below Token-MIS at $28.85$), yielding the strongest aggregate performance across tasks. In contrast, although the accuracy of token and sequence level MIS does not collapse, their training-inference KL or policy-update KL blow up during the training.
ALP enhances training stability.
Figure 4 further show that Seq-ALP exhibits consistently well-behaved optimization dynamics. In particular, Seq-ALP maintains controlled train–inference KL without the pronounced excursions observed in Token-MIS, and avoids the sharp spikes in policy-update KL that characterize Seq-MIS. In addition, Seq-ALP maintains moderately higher entropy (supporting exploration) while avoiding the entropy blow-up in Seq-Bypass, indicating that ALP stabilizes both policy drift and update magnitude during training.
@@
Perturbation promotes exploration efficiency.
Finally, the Pass@$k$ curves on AIME24 and 25 provide direct evidence that latent-space perturbations translate into more effective exploration (Figure 5). Across moderate-to-large rollout budgets $k$ (1625616–256), ALP consistently attains the highest Pass@$k$, with the advantage becoming more pronounced as $k$ increases, suggesting that perturbing hidden states helps generate a more diverse set of candidate solution trajectories and thus improves the probability of sampling a correct reasoning path under fixed decoding temperature.
@@ 6.1 Perturbation targets
Figure 6: Perturbation-layer ablation in the single-turn setting. The results reveal a clear performance hierarchy: All layers >> Partial layers >> Logits only; Bottom layers >> Top layers.
@@
7 Conclusions and Future Work
We introduced Adaptive Layerwise Perturbation (ALP), a simple and unified approach for stabilizing off-policy RL in large language models. ALP injects learnable perturbations into transformer hidden states when computing the updated policy, and optimizes a single importance ratio between the perturbed training distribution and the rollout distribution. This perspective addresses both policy staleness (multiple mini-batch updates per rollout) and training–inference mismatch within a single formulation, while avoiding the complexity of separately truncating multiple ratios. Empirically, ALP consistently improves stability and performance on both single-turn and multi-turn agentic reasoning tasks. Ablations confirm that perturbing hidden states is essential, with perturbing all layers performing best overall and lower-layer perturbations typically preferred when only a subset is perturbed.
Several directions remain promising. First, we will evaluate ALP in fully asynchronous RL systems where a single trajectory is generated under several sequential policies; this setting exacerbates off-policy effects and provides a stringent test of ALP’s ability to control heavy-tailed ratios. Second, we will extend ALP to Mixture-of-Experts (MoE) models, where routing decisions and expert imbalance can amplify distribution shift and instability; perturbations may need to be coupled with routing-aware constraints or expert-specific scales. More broadly, it would be valuable to develop principled strategies for allocating perturbation across layers and experts, and to understand how perturbation interacts with truncation-based methods, quantization, and other systems-level sources of mismatch.
Appendix A Technical Lemmas
@@
Appendix B Formal Statement and Proof for Theorem 1
@@
Condition 1 requires that the squared $\ell_{2}$ norm of the posterior mean of the perturbation is bounded by $\mathcal{O}(d\sigma^{2})$, where $d$ is the input dimension and $\sigma^{2}$ is the variance of the Gaussian prior $\delta\sim\mathcal{N}(0,\sigma^{2}I_{d})$. This scaling is natural and aligns with standard concentration behavior in high-dimensional Gaussian settings. Condition 2 ensures that the perturbed policy $\widetilde{\pi}_{\theta}$ is uniformly dominated by the original policy $\pi\theta$ up to a constant factor $\alpha$. Intuitively, this condition guarantees that the injected perturbation is well controlled and does not overly distort the action distribution, which could otherwise lead to instability during training. In the limiting case $\delta\to 0$, the condition is trivially satisfied with $\alpha=1$.
@@
Appendix C Proof for Theorem 2
To start with, we denote $b:=max_{x}\mathcal{I}(x,\theta)$, and there exists $c\in(0,2)$ such that
@@ Appendix C Proof for Theorem 2
where the last inequality is from the fact that the measure of set satisfying $\mathcal{I}(x,\theta)\geq c/2\cdot b$ equals to $C$, and for a Gaussian distribution, the probability on such set is smaller than the probability on a zero-center ball with volume $C$. Combing this fact with the assumption in Equation (8), we could derive the last inequality.
@@
Appendix D Additional Expeimental Details

### 29 Apr 2026 v2

@@ 3.1.1 Experimental settings
Dataset. For single-turn reasoning, we study math tasks without tool use. The training set merges the math subset of Guru RL-92k [^2] https://huggingface.co/datasets/LLM360/guru-RL-92k(Cheng et al., 2025) and a 7575K OpenR1333https://huggingface.co/datasets/weqweasdas/from_default_filtered_openr1_with_scores(Hugging Face, 2025) subset, with correctness verified by Math-Verify (Kydlíček, ). Additional preprocessing details are deferred to Appendix E.
Training. We implement all methods in verl (Sheng et al., 2024), following its default setup where possible. We use AdamW with learning rate $1\times 10^{-6}$ and the clip-higher trick (Yu et al., 2025) with asymmetric clipping $(1-\epsilon_{l},1+\epsilon_{h})$. The models are trained from Qwen2.5-Math-1.5B-base (Yang et al., 2024); each iteration samples $512$ prompts, rolls out $n=8$ responses per prompt at temperature $1.0$, and performs $16$ policy updates. As provided in Table 5 in the appendix, token/seq denotes whether the ratio is aggregated at the token level or sequence level. The definitions of these notations are provided in Sec. 2.1. Unless otherwise specified, the main single-turn experiments use Qwen2.5-Math-1.5B-base; we additionally report a validation run on Qwen3-4B (Yang et al., 2025) to test whether observed trend transfers to a different backbone.
Benchmarks. We evaluate on Math500 (Hendrycks et al., 2021), Minerva Math (Lewkowycz et al., 2022), Olympiad Bench (He et al., 2024), AIME2024444https://huggingface.co/datasets/math-ai/aime24, and AIME2025555https://huggingface.co/datasets/math-ai/aime25. We report average@32 at temperature $1.0$ with a generation limit of $4096$ tokens. Baselines include Seq-Bypass, Seq-MIS, Token-MIS, and vanilla GRPO; Table 5 summarizes how each method forms and clips the importance ratios.
@@ 3.1.2 Main Results
Figure 3: Single-turn training dynamics under training-inference mismatch: reward, gradient norm, entropy, and rollout-training KL over policy-update steps. Reward is smoothed with a 10-step moving average.
Method Math500 Minerva Math Olympiad Bench AIME24 AIME25 Average
GRPO 94.49 56.96 68.28 60.29 51.17 66.24
Token-MIS 94.28 53.66 66.36 62.35 50.00 65.33
Seq-Bypass 93.70 52.00 64.21 63.80 46.88 64.12
Seq-ALP 95.18 57.44 69.49 64.32 51.17 67.52
Table 2: Single-turn validation on Qwen3-4B under the same evaluation protocol. Seq-ALP achieves the best average score across benchmarks.
ALP-type delivers the top performance. Under single-turn evaluation in Table 1, Token-ALP attains the best overall average score ($37.87$), with Seq-ALP second ($36.83$). Token-MIS is the strongest baseline, while Seq-MIS, GRPO, and Seq-Bypass perform worse overall. To test whether the benefit of perturbation is specific to Qwen2.5-Math-1.5B, we further repeat the single-turn experiment on Qwen3-4B. Table 2 shows that Seq-ALP achieves the best overall average score ($67.52$), outperforming GRPO ($66.24$), Token-MIS ($65.33$), and Seq-Bypass ($64.12$). This suggests that the gain from ALP is not tied to one particular base model and persists at a larger model scale.
ALP consistently improves training stability. Figure 3 shows that both token-level and sequence-level ALP remain stable throughout training: gradient norms stay controlled, entropy remains in a reasonable range, and rollout-training KL stays near zero. By contrast, GRPO and Seq-MIS exhibit late-stage spikes in gradient norm and KL together with reward collapse, while Seq-Bypass and Token-MIS are more stable but still less competitive than ALP overall.
@@ 4.1 Experimental settings
We study multi-turn Tool-Integrated Reasoning (TIR) with a Python interpreter. We adapt the SimpleTIR codebase (Xue et al., 2025), mask void turns in the loss, and train on data merged from SimpleRL (Zeng et al., 2025), Deepscaler (Luo et al., 2025), and rStar2-Agent (Shang et al., 2025). ALP follows the zero-RL setting and starts from Qwen2.5-7B-base. The rollout batch size is $512$, the mini-update size is $32$, the maximum response length is 1616K, and each episode allows up to $5$ code-execution turns. We use the same baselines and evaluation benchmarks as in the single-turn setting.
@@ 4.2 Main Results
Figure 4: Multi-turn training dynamics. Seq-ALP maintains higher but controlled entropy and stable convergence in both KL metrics, while Token-MIS is unstable in train-inference KL and Seq-MIS shows spikes in policy-update KL.
Figure 5: Pass@$k$ on AIME 2024 and AIME 2025 for TIR. ALP achieves the highest accuracy over rollout budgets $k=16\sim 256$, indicating improved exploration efficiency and solution diversity.
ALP has the best performance. On multi-turn TIR benchmarks in Table 3, Seq-ALP achieves the highest overall average accuracy ($50.53$), outperforming Token-MIS, Seq-MIS, and Seq-Bypass. The gains are broad-based across benchmarks, with only AIME25 slightly favoring Token-MIS.
ALP enhances training stability. Figure 4 shows that Seq-ALP maintains controlled train-inference KL and preserves higher but stable entropy. While the reward for MIS does not collapse, their KL divergence still collapse and the additionally masking easily results in early plateau.
@@
Perturbation promotes exploration efficiency. Finally, the Pass@$k$ curves on AIME24 and 25 provide direct evidence that latent-space perturbations translate into more effective exploration (Figure 5). Across moderate-to-large rollout budgets $k$ (1616256256), ALP consistently attains the highest Pass@$k$, with the advantage becoming more pronounced as $k$ increases, suggesting that perturbing hidden states helps generate a more diverse set of candidate solution trajectories and thus improves the probability of sampling a correct reasoning path under fixed decoding temperature.
@@ 5.1 Perturbation targets
Figure 6: Perturbation-layer ablation in the single-turn setting. The results reveal a clear performance hierarchy: All layers >> Partial layers >> Logits only; Early layers >> Late layers.
@@
6 Conclusions
We introduced Adaptive Layerwise Perturbation (ALP), a unified approach for stabilizing off-policy RL in large language models. ALP injects learnable hidden-state perturbations when computing the updated policy and optimizes a single importance ratio against the rollout distribution, thereby addressing policy staleness and training-inference mismatch within one formulation. Our analysis suggests that perturbation helps both by tightening the train-infer discrepancy and by smoothing sharp local optimization geometry. Empirically, across single-turn and multi-turn reasoning tasks, ALP improves both training stability and final performance, while ablations show that hidden-state perturbation, especially across all layers, is most effective. Promising future directions include fully asynchronous RL and Mixture-of-Experts settings, where off-policy effects and systems mismatch can be even more severe.
Appendix A Additional Related Works
Off-policy
Recent systems efforts aim to eliminate training-inference mismatch by enforcing bitwise-consistent inference, enabling closer-to on-policy RL in practice (He and Lab, 2025) or using higher precisions (Qi et al., 2025). These directions are promising, but maintaining strict bitwise consistency is computationally inefficient, and can be challenging under common production constraints such as changing execution engines, optional rollout quantization, fully asynchronous rollouts, and rapidly evolving kernels. Our work is complementary: rather than assuming mismatch can be removed everywhere, we propose an algorithmic correction that remains robust when such mismatch persists, providing a unified importance-ratio formulation that tolerates heterogeneous sources of off-policy deviation. Although those methods are effective, they only focus on a specific problem and current RL training needs to combine the techniques and tune the parameters separately. The algorithm becomes even more complicated in fully-asynchronous settings (Fu et al., 2025) since there are three importance ratios. Hence, our work develops a unified importance ratio instead of dividing the importance ratio into several parts to solve general off-policy problems.
Appendix B Technical Lemmas
@@
Appendix C Formal Statement and Proof for Theorem 1
@@
Condition 1 requires that the squared $\ell_{2}$ norm of the posterior mean of the perturbation is bounded by $\mathcal{O}(d\sigma^{2})$, where $d$ is the input dimension and $\sigma^{2}$ is the variance of the Gaussian prior $\delta\sim\mathcal{N}(0,\sigma^{2}I_{d})$. This scaling is natural and aligns with standard concentration behavior in high-dimensional Gaussian settings. Condition 2 ensures that the perturbed policy $\widetilde{\pi}_{\theta}$ is uniformly dominated by the original policy $\pi_{\theta}$ up to a constant factor $\alpha$. Intuitively, this condition guarantees that the injected perturbation is well controlled and does not overly distort the action distribution, which could otherwise lead to instability during training. In the limiting case $\delta\to 0$, the condition is trivially satisfied with $\alpha=1$.
@@
Appendix D Proof for Theorem 2
Theorem 4.
Define $\mathcal{I}(x,\theta)=\|\nabla^{2}_{\theta}J(\theta)\|_{2},\quad\widetilde{\mathcal{I}}(x,\theta)=\|\nabla^{2}_{\theta}J_{\mathrm{ALP}}(\theta,\sigma)\|_{2}$. If there exists a smallest constant $0<c<1$ such that
To start with, we denote $b:=\max_{x}\mathcal{I}(x,\theta)$, and there exists $c\in(0,2)$ such that
@@ Appendix D Proof for Theorem 2
where the last inequality is from the fact that the measure of the set satisfying $\mathcal{I}(x,\theta)\geq c/2\cdot b$ equals $C$, and for a Gaussian distribution, the probability on such a set is smaller than the probability on a zero-center ball with volume $C$. Combining this fact with the assumption in Equation (7), we derive the last inequality.
@@
Appendix E Additional Experimental Details
Single-turn details.
@@ Appendix E Additional Experimental Details
We further provide the clipping or masking details for the baselines in the following table.
GRPO token None None
Token-MIS token seq None
Seq-MIS seq seq None
Seq-Bypass None None seq
Table 5: Settings of parameters of importance ratios
@@
Multi-turn details.
For multi-turn experiments, we use Tool-Integrated Reasoning (TIR) with a Python interpreter, allowing the model to execute code for intermediate verification and symbolic computation. We adapt the SimpleTIR codebase and mask void turns in the loss, where void turns are turns without tool calls and without final answers. The training data merges Math3-5 examples from SimpleRL, Deepscaler, and rStar2-Agent. Training starts from Qwen2.5-7B-base under the zero-RL setting, with rollout batch size $512$, mini-update size $32$, maximum response length 1616K, and at most $5$ code-execution turns per episode.